Construct a truth table for each of these compound propositions.

(i) (p → q) ↔ (¬q → ¬p) (16 marks)

ii) p ⊕ (p ∨ q) (8 marks)


(i) Determine by using truth tables if (p ∧ q) → p is a tautology, contradiction or a contingency. Give reasons for your answer. (6 marks)

(ii) Show that ¬(p ⊕ q) and p ↔ q are logically equivalent. (6 marks)


1
Expert's answer
2021-12-08T15:49:03-0500

Let us construct a truth table for each of these compound propositions.


(i) "(p \u2192 q) \u2194 (\u00acq \u2192 \u00acp)"


"\begin{array}{||c|c||c|c|c|c|c||} 
\hline\hline 
p & q & p → q & \neg q & \neg p & ¬q → ¬p & (p → q) ↔ (¬q → ¬p) \\ 
\hline\hline 
0 & 0 & 1 & 1 & 1 & 1 & 1\\ 
\hline 
0 & 1 & 1 & 0 & 1 & 1 & 1 \\
 \hline 
1 & 0 & 0 & 1 & 0 & 0 & 1\\ 
\hline 
1 & 1 & 1 & 0 & 0 & 1 & 1\\ 
\hline\hline 
\end{array}"


ii) "p \u2295 (p \u2228 q)"


"\begin{array}{||c|c||c|c|c|c|c||} 
\hline\hline 
p & q & p ∨ q & p ⊕ (p ∨ q) \\ 
\hline\hline 
0 & 0 & 0 & 0\\ 
\hline 
0 & 1 & 1 & 1 \\
 \hline 
1 & 0 & 1 & 0\\ 
\hline 
1 & 1 & 1 & 0 \\ 
\hline\hline 
\end{array}"


(i) Let us determine by using truth tables if "(p \u2227 q) \u2192 p" is a tautology, contradiction or a contingency.


"\begin{array}{||c|c||c|c||} 
\hline\hline 
p & q &p ∧ q & (p ∧ q) → p \\ 
\hline\hline 
0 & 0 & 0 & 1\\ 
\hline 
0 & 1 & 0 & 1
 \\ \hline 
1 & 0 & 0 & 1\\ 
\hline 
1 & 1 & 1 & 1\\ 
\hline\hline 
\end{array}"


Since the last column contains only 1, we conclude that this formula is a tautology. Therefore, this formula neither a contradiction, nor a contingency.


(ii) Let us show that "\u00ac(p \u2295 q)" and "p \u2194 q" are logically equivalent using the truth table.


"\begin{array}{||c|c||c|c|c||} 
\hline\hline 
p & q & p ⊕ q & ¬(p ⊕ q) & p ↔ q \\ 
\hline\hline 
0 & 0 & 0 & 1 & 1\\ 
\hline 
0 & 1 & 1 & 0 & 0
 \\ \hline 
1 & 0 & 1 & 0 & 0\\ 
\hline 
1 & 1 & 0 & 1 &1\\ 
\hline\hline 
\end{array}"


Since the last two columns are coinside, the formulas "\u00ac(p \u2295 q)" and "p \u2194 q" are logically equivalent.